OpenStudy (anonymous):

find the limit of sqrt[(x)(x+c)] - x where c is a constant as x approaches positive infinity

OpenStudy (freckles):

\[\lim_{x \rightarrow \infty}(\sqrt{x^2+xc}-x)\] is this correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

looks like negative infinity

OpenStudy (freckles):

\[\lim_{x \rightarrow \infty}\frac{\sqrt{x^2+xc}-x}{1} \cdot \frac{\sqrt{x^2+xc}+x}{\sqrt{x^2+xc}+x}\] Try multiply by top's conjugate on bottom and top like i have here

OpenStudy (freckles):

then look at the bottom to see what to divide top and bottom by

OpenStudy (anonymous):

ok give me a minute to compute :)

OpenStudy (anonymous):

i got -x^2/1 lol. my calc class is on extrema though no on infinite limits yet

OpenStudy (anonymous):

so far I have \[\frac{ x^{2}+xc-x }{ \sqrt{x ^{2}+xc}+x} \]

OpenStudy (freckles):

x*x=x^2 \[\frac{ x^{2}+xc-x^2 }{ \sqrt{x ^{2}+xc}+x}\]

OpenStudy (anonymous):

OH!

OpenStudy (freckles):

\[\lim_{x \rightarrow \infty} \frac{xc}{\sqrt{x^2+xc}+x}\]

OpenStudy (anonymous):

so where should I go from here?

OpenStudy (freckles):

so you want to divide both top and bottom by the highest degree on bottom but you have a sqrt on bottom so divide both top and bottom by sqrt(x^2) and sqrt(x^2)=x if x>0 sqrt(x^2)=-x if x<0 we are approaching positive infinity so we are looking at using sqrt(x^2)=x

OpenStudy (anonymous):